OpenStudy (anonymous):

how do find the equation of the tangent and the equation of the secant of y=x^2 at x=-2 ?

OpenStudy (anonymous):

for the secant line you would need two points for the line tangent at \((-2,4)\) you need a slope, which is given by the derivative evaluated at \(x=-2\)

OpenStudy (anonymous):

ive gotten that the slope is -4

OpenStudy (anonymous):

the derivative of \(f(x)=x^2\) is \(f'(x)=2x\) and \(f'(-2)=-4\) meaning the slope of your line is \(-4\)

OpenStudy (anonymous):

ok good last step is to use the point slope formula since you have the slope is \(-4\) and the point is \((-2,4)\)

OpenStudy (anonymous):

this is for the tangent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for a "secant" line you are going to need two points on the curve, not one

OpenStudy (anonymous):

ok. i got y = -4x - 4. now for the secant, how do i get two points?

OpenStudy (anonymous):

the question is not well posed you cannot be asked to find the equation of "the" secant line at only one point. you need to be told two points. is that exactly how the question reads?

OpenStudy (anonymous):

hmm. it just says: "at the indicated point find a) the slope of the curve b) the equation of the tangent, and c) an equation of the normal wait... so then im guessing "secant" and "normal" arent the same thing... haha.

OpenStudy (anonymous):

ooooooooooooooh no

OpenStudy (anonymous):

you want the "normal line" that means the line perpendicular to the curve at that point

OpenStudy (anonymous):

soooo reciprocal of -4 and find b

OpenStudy (anonymous):

but you have done the main part of the work already slope of tangent line is \(-4\) so slope of perpendicular line is \(\frac{1}{4}\)

OpenStudy (anonymous):

careful, not the reciprocal, the negative reciprocal point is of course the same so it is just another application of the point slope formula

OpenStudy (anonymous):

ahhh yes thats what i meant. i got y=(1/4)x + (9/2) thanks! (:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh and yes, that is right